# The SWE-Bench Illusion: When State-of-the-Art LLMs Remember Instead of Reason

### v2

@@ 3.3 Metrics
In this section, we details our choice of metrics used for our experiments.
@@
To address concerns that models might succeed through superficial pattern matching rather than genuine understanding, we introduce filtered accuracy as a control metric. This measures model performance exclusively on issue instances where problem descriptions do not explicitly contain file paths or import statements, identified through simple heuristics. By removing cases where path-like strings appear in issue descriptions, filtered accuracy isolates performance that depends purely on repository knowledge and problem-solving reasoning rather than surface-level string matching. This metric directly addresses our research question of whether models demonstrate genuine coding understanding or rely on memorized patterns.
@@
5-gram Consecutive Accuracy (5-gram Acc.)
To quantify the degree of verbatim memorization in model outputs, we compute 5-gram consecutive accuracy, which measures the proportion of predicted 5-grams (consecutive sequences of 5 tokens) that exactly match those found in the ground truth. This metric is calculated as the ratio of matching 5-grams to the total number of 5-grams in the prediction, with frequency-aware matching to prevent inflation from repeated phrases. Given that random 5-gram matches are statistically improbable, high scores on this metric provide strong evidence of memorization rather than genuine code generation. This allows us to distinguish models that have memorized data patterns in the training samples.
@@ 3.4 Models
We evaluated a diverse set of ten state-of-the-art LLMs from both OpenAI and Anthropic to measure their ability on file path identification task. From OpenAI, we included snapshot variants of GPT‑4o (gpt‑4o‑2024‑08, and gpt‑4o‑2024‑05), alongside GPT‑4.1 (gpt-4.1-2025-04-14), the reasoning-focused OpenAI o3 (o3-2025-04-16) and its lightweight counterpart o3‑mini (o3-mini-2025-01-31), as well as o4‑mini (o4-mini-2025-04-16), the compact successor OpenAI (2025a). From Anthropic, our suite comprised the Claude family: Claude 3.5 Sonnet (June 20, 2024), Claude 3.7 Sonnet (February 24, 2025), Claude 4.0 Sonnet, and Claude 4.0 Opus, both officially released on May 22, 2025.
For consistency, OpenAI’s chat models and Claude models were configured with a maximum generation length of 2,048 completion tokens. For the reasoning-centric “o” family from OpenAI (i.e., o3, o3‑mini, o4‑mini), we increased the limit to 4,096 tokens. All experiments used the respective provider APIs with default sampling settings to reflect typical usage.
@@ 4.1 File Path Identification Accuracy
4.1.1 Instance-Specific memorization
@@
The first evidence pattern emerges from performance differences within the SWE-Bench repository ecosystem. Models consistently achieve their highest accuracy on the curated SWE-Bench-Verified subset (60-76%), with a noticeable decline on the broader Full-SWE-Bench set (57-71%) and a further drop on newly collected SWE-Repo tasks (50-68%).
@@
This graduated performance decay suggests that models’ knowledge is not uniform even across the same codebases. The superior performance on the ”Verified” set, which is the most commonly used for evaluation and reporting, points toward instance-specific memorization, where models have been disproportionately exposed to or optimized for these specific, widely-circulated benchmark problems.
4.1.2 Repository-Bias memorization
@@
The second, more dramatic pattern appears when comparing performance between SWE-Bench and external benchmarks. While this performance drop is partly compounded by confounding variables, such as the significantly shorter issue descriptions in RefactorBench (Table 1), the data still points overwhelmingly to repository-level bias.
@@
The most compelling evidence comes from the Outside-Repo Tasks. These tasks were sourced from popular repositories (pandas, PyTorch) that are almost certainly well-represented in training data. If models developed a general capability for popular Python projects, we would expect strong performance here. Instead, accuracy is uniformly lower (below 53%) than those in SWE-Bench Verified.
@@
The stark contrast between high performance on the SWE-Bench repositories and poor performance on these other popular repositories refutes the idea that models are simply good at ”popular projects.” It strongly supports our hypothesis of repository-bias memorization, where models have overfit to the specific architectural patterns and problem distributions of the dozen repositories in the SWE-Bench collection, failing to develop transferable skills.
4.1.3 Model-Specific Contamination Patterns
@@
The results reveal differences in contamination patterns across model families and vendors:
@@
OpenAI chat Models: OpenAI’s chat model variants show inconsistent performance patterns across benchmarks, with no stable ranking order. The performance ranking of these models varies significantly between different benchmark types. The only consistent pattern is GPT-4o-2024-08-06, which maintains mediocre performance across most evaluations.
@@
OpenAI reasoning Models: While o3 demonstrates higher accuracy than o3-mini and o4-mini due to its enhanced capabilities, the performance patterns of o3-mini and o4-mini follow similar trends to the chat models. We cannot identify a consistently stronger model within this family for the path prediction task across different benchmarks.
@@
Anthropic Models: Claude models exhibit more consistent performance patterns, with newer model generations generally showing higher path prediction accuracy. However, Claude Sonnet 4 shows notably lower performance than other Claude family models on RefactorBench and Extra-Repo tasks, which represents an exception to the otherwise consistent generational improvement pattern. Overall, Claude Sonnet 4 and Opus 4 maintain relatively higher performance across most benchmarks compared to other Claude models.
@@
Cross-Vendor Consistency: A key finding from our evaluation of ten different LLMs is the remarkable consistency of these patterns across different model families and vendors. The fact that all models, regardless of vendor, exhibit the same performance hierarchy (SWE-bench Verified >>> Full-SWE-Bench >>> SWE-Repo Tasks >>> SWE-bench C#, RefactorBench, and Outside-Repo Tasks) indicates the performance disparities reflect systematic exposure patterns in training data rather than isolated vendor-specific issues.
4.1.4 Implications for Benchmark Validity
@@
The results provide strong empirical evidence for the two forms of memorization hypothesized in our introduction. The performance degradation within the SWE-Bench ecosystem points to instance-specific memorization, while the more significant performance collapse on external repositories indicates repository-bias memorization. The dual presence of these phenomena suggests that high scores on SWE-Bench are not purely indicative of generalizable coding ability. Instead, they are likely inflated by these confounding factors, reflecting a specialized knowledge of the benchmark’s specific instances and repository patterns rather than transferable software engineering skills.
@@
SWE-Repo Tasks 100 (29.1%) 244 (70.9%) 344
SWE-bench C# 11 (14.7%) 64 (83.1%) 77
@@
Outside-Repo Tasks 26 (16.5%) 132 (83.5%) 158
Table 2: Dataset splits with corresponding proportions (as percentages) for the instances that contains ground truth path information and those do not.
@@ 4.2 File Path Identification Accuracy on Filtered Instances
First, models still achieve their highest accuracy on SWE-Bench Verified instances, followed by Full SWE-Bench and SWE-Repo tasks confirming a pattern of instance-specific memorization that is not dependent on clues from the prompt. Second, the sharp performance drop on external repositories also persists, providing definitive evidence for repository-bias memorization.
The persistence of these patterns in the filtered data is our most critical finding. It shows that the models’ ability to identify correct files is not due to simple text matching but is based on memorized knowledge, which includes instance-specific patterns and repository-biased patterns. This suggests that current benchmark scores are inflated by these memorization effects and do not purely reflect a model’s transferable software engineering skills.
Figure 6: Comparison of models’ consecutive 5-gram accuracy on function reproduction across benchmarks. Red ticks indicate improved verbatim similarity with the ground truth code, while blue ticks indicate a decrease.
@@ 4.3 Function Reproduction Result
We analyze model performance on the function reproduction task in this subsection. As shown in Figure 6, models achieve up to 35.0% consecutive 5-gram accuracy reproducing ground truth functions (the lower plot) from SWE-Bench Verified compared to up to 18.2% on instances outside the SWE-Bench ecosystem666 Due to limited resources, we only conducted experiments for o3-mini for the OpenAI reasoning models. . With limited context provided, only function names and issue descriptions, models lack sufficient information to systematically derive correct implementations. High consecutive 5-gram accuracy under these constraints requires reproducing exact textual sequences, indicating reliance on memorized training data rather than reasoning.
@@
: Performance reveals a clear hierarchy that mirrors benchmark curation processes: SWE Verified (35.0%) >>> SWE Full (28.9%) >>> SWE Extra (18.3%). Critically, SWE Extra performance drops to the same level as external benchmarks (RefactorBench: 18.2%, Outside Tasks: 17.7%), despite originating from identical repositories as SWE Verified. This pattern indicates that models have specifically memorized the curated, canonical solutions that human evaluators deemed high-quality, rather than developing general familiarity with repository codebases.
These results demonstrate that SWE-Bench Verified performance substantially reflects memorization of training sequences rather than coding competence. The inability to reproduce functions from less-curated instances within the same repositories, combined with performance equivalence between SWE Extra and external benchmarks, reveals that apparent programming ability is largely an artifact of exposure to the specific, high-quality solutions selected for benchmark inclusion.
@@ 5.1 Coding Benchmarks and Agents
Various coding benchmarks have been proposed to evaluate LLMs and LLM agents’ coding abilities, including SWE-Bench, BigCodeBench, SWE-Gym, EvoEval Jimenez et al. (2024); Zhuo et al. (2024); Xia et al. (2024). Among them, SWE-Bench has emerged as one of the most popular evaluation method for LLM’s coding abilities (Jimenez et al., 2024). The benchmark consists of real GitHub issues and their corresponding patches, providing a seemingly realistic evaluation setting. However, the public nature of GitHub data raises concerns about potential exposure during training. Recent agent systems have shown dramatic improvements on SWE-Bench (Yang et al., 2024; Wang et al., 2024), but the extent to which these improvements reflect genuine problem-solving versus pattern matching remains unclear.
@@ 5.2 Benchmark Contamination
The issue of benchmark contamination, where models are inadvertently trained on evaluation data, is a well-documented problem that can lead to inflated performance metrics without genuine understanding Zhou et al. (2023). To specifically combat this in the coding domain, prior work has developed two primary strategies: task mutation and metric-based probing.
Task mutation approaches aim to detect memorization by altering existing benchmark problems. For example, EvoEval Xia et al. (2024) uses LLMs to create semantic-altering transformations of HumanEval Chen et al. (2021) problems, while other methods Chen et al. (2025) use AST-based mutations to measure how performance changes with structural code modifications. Similarly, TaskEval Tambon et al. (2025) measures prompt sensitivity by creating numerous variations of the same task. These methods are effective at revealing overfitting on specific problem formulations.
Metric-based probing approaches, in contrast, analyze the model’s outputs without changing the task itself. These techniques Li (2023); Ramos et al. (2025) include using perplexity analysis or other statistical measures like Negative Log Likelihood (NLL) and n-gram accuracy to infer if a model has likely seen a specific bug or solution before. However, perplexity and NLL approaches cannot be applied to commercial models, for we do not have access to the model’s hidden states. N-gram similarity, while applicable, can be a noisy indicator for complex code patches; any functionally correct solution will inevitably share significant textual overlap with the ground truth, making it difficult to distinguish genuine problem-solving from verbatim memorization.
While existing mutation and metric-based methods are effective for single-function code generation, they face limitations when applied to complex, multi-file software engineering benchmarks like SWE-Bench. Task mutation approaches require generating semantically equivalent variations of real-world GitHub issues, a significantly more challenging task than mutating isolated coding problems. Metric-based approaches using perplexity or NLL cannot be applied to commercial models due to lack of access to hidden states, while n-gram similarity becomes noisy for complex patches where any functionally correct solution will share substantial textual overlap with ground truth.
Our work introduces a novel differential testing framework that addresses these limitations by using comparative performance analysis across related benchmarks as memorization proxies. Rather than requiring ground truth knowledge of training data exposure, we systematically compare model performance across benchmark variants and external evaluations to detect suspicious performance patterns. This differential testing methodology provides a systematic, reusable framework for detecting memorization in any coding benchmark without requiring access to training data or model internals.

### v3

@@ 3.3 Metrics
In this section, we discuss metrics used in our experiments.
@@
To address concerns that models might succeed through superficial pattern matching rather than genuine understanding, we introduce filtered accuracy as a control metric. This measures performance exclusively on instances where problem descriptions do not explicitly contain file paths or import statements, identified through simple heuristics. By removing cases where path-like strings appear in issue descriptions, filtered accuracy isolates performance that depends purely on repository knowledge and problem-solving reasoning rather than surface-level string matching. This metric directly addresses our research question of whether models demonstrate genuine coding understanding or rely on memorized patterns.
@@
5-gram Consecutive Overlap (5-gram Overlap.)
To quantify the degree of verbatim memorization in model outputs, we compute 5-gram overlap ratio, which measures the proportion of predicted 5-grams (consecutive sequences of 5 tokens) that exactly match those found in the GT code. This metric is calculated as the ratio of matching 5-grams to the total number of 5-grams in the prediction, with frequency-aware matching that respects the occurrence count of each n-gram in the ground truth to prevent inflation from repeated phrases. This metric captures the extent to which models regurgitate exact token sequences. When models show high overlap on some benchmarks but not others, this differential pattern can indicate memorization of specific training data rather than consistent coding ability.
$\text{N-gram Overlap (\%)}=\frac{\text{Number of matched n-grams}}{\text{Total generated n-grams}}\times 100$ (2)
Instance-Level Verbatim Match Percentage
For the prefix completion task, an instance is marked as compromised if the model generates at least one code hunk that exactly matches the corresponding GT implementation. A high instance-level verbatim match percentage suggests that many test instances may be untrustworthy, as portions of the output are memorized by the model. We report the percentage of compromised instances across the entire benchmark.
@@ 3.4 Models
We evaluated a diverse set of ten state-of-the-art LLMs from both OpenAI and Anthropic to measure their ability on file path identification task. From OpenAI, we included snapshot variants of GPT‑4o (gpt‑4o‑2024‑08, and gpt‑4o‑2024‑05), alongside GPT‑4.1 (gpt-4.1-2025-04-14), reasoning-focused OpenAI o3 (o3-2025-04-16) and its lightweight counterpart o3‑mini (o3-mini-2025-01-31), as well as o4‑mini (o4-mini-2025-04-16), the compact successor (OpenAI 2025a). From Anthropic, our suite comprised the Claude family: 3.5 Sonnet (June 20, 2024), 3.7 Sonnet (February 24, 2025), 4.0 Sonnet, and 4.0 Opus, both officially released on May 22, 2025.
For consistency, OpenAI’s chat models and Claude models were configured with a maximum generation length of 2,048 completion tokens. For the reasoning-centric “o” family (i.e., o3, o3‑mini, o4‑mini), we increased the limit to 4,096 tokens. All experiments used the respective provider APIs with default sampling settings to reflect typical usage.
@@ 4.1 File Path Identification Accuracy
Instance-Specific memorization
@@
The first evidence pattern emerges from performance differences within the SWE-Bench repo ecosystem. Models consistently achieve their highest accuracy on the curated SWE-Bench-Verified subset (60-76%), with a noticeable decline on the broader Full-SWE-Bench set (57-71%) and a further drop on newly collected SWE-Repo tasks (50-68%).
@@
This graduated performance decay suggests that models’ knowledge is not uniform even across the same codebases. The superior performance on the “Verified” set, which is the most commonly used for evaluation and reporting, points toward instance-specific memorization, where models have been disproportionately exposed to or optimized for these specific, widely-circulated benchmark problems.
Repository-Bias memorization
@@
The second, more dramatic pattern appears when comparing performance between SWE-Bench and external benchmarks. While this performance drop is partly compounded by confounding variables, such as the significantly shorter issue descriptions in RefactorBench (Table 1), the data still points overwhelmingly to repo-level bias.
@@
The most compelling evidence comes from the Outside-Repo Tasks. These tasks were sourced from popular repositories that tend to be well-represented in training data. If models developed a general capability for popular Python projects, we would expect strong performance here. Instead, accuracy is uniformly lower (<<53%) than that in SWE-Bench Verified.
@@
This stark contrast between high performance on SWE-Bench repositories and poor performance on equally popular repositories reveals that model success depends on more than just repository popularity. It thus supports our hypothesis of repository-bias memorization, where models have overfit to the specific architectural patterns and problem distributions of the twelve repositories in the SWE-Bench collection, failing to develop transferable skills.
Model-Specific Contamination Patterns
@@
The results reveal differences in contamination patterns across model families/vendors:
@@
OpenAI chat Models: OpenAI’s chat models show inconsistent performance patterns across benchmarks, with no stable ranking order. The performance ranking of these models varies significantly between different benchmark types. The only consistent pattern is GPT-4o-2024-08-06, which shows mediocre performance across most experiments.
@@
OpenAI reasoning Models: While o3 demonstrates higher accuracy than o3-mini and o4-mini due to its enhanced capabilities or higher degree of overfitting, the performance patterns of o3-mini and o4-mini follow similar and inconsistent trends to the chat models.
@@
Anthropic Models: Claude models exhibit more consistent performance patterns, with newer model generations showing higher path prediction accuracy. However, Claude Sonnet 4 shows notably lower performance than other Claude family models on RefactorBench, representing an exception to the otherwise consistent generational improvement pattern. Overall, Sonnet 4 and Opus 4 maintain relatively higher performance across most benchmarks compared to other Claude models.
@@
Cross-Vendor Consistency: A key finding from our evaluation of ten different LLMs is a remarkable consistency of these patterns across different model families and vendors. The fact that all models, regardless of vendor, exhibit the same performance hierarchy (SWE-bench Verified >> Full-SWE-Bench >> SWE-Repo Tasks >> SWE-bench C#, RefactorBench, and Outside-Repo Tasks) indicates the performance disparities reflect systematic exposure patterns in training data rather than issues isolated to a vendor.
Implications for Benchmark Validity
@@
The results provide strong empirical evidence for the two forms of memorization hypothesized in our introduction. The performance degradation within the SWE-Bench style benchmarks points to instance-specific memorization, while the more significant performance degradation on external repositories indicates repository-bias memorization. The presence of these two types of memorization suggests that high scores on SWE-Bench are not purely indicative of generalizable coding abilities for these models. Instead, they are likely inflated by these confounding factors. This shows that models possess specialized knowledge of specific tasks in SWE-Bench Verified and repository patterns, rather than or possibly in addition to transferable software engineering skills.
@@
SWE-Repo Tasks 57 (26.3%) 160 (73.7%) 217
Outside-Repo Tasks 63 (25.7%) 182 (74.3%) 245
@@
SWE-bench C# 11 (14.7%) 64 (83.1%) 75
Table 2: Dataset splits with corresponding proportions (as percentages) for the instances that contain ground truth path information and those that don’t.
@@ 4.2 File Path Identification Accuracy on Filtered Instances
First, models still achieve the highest accuracy on SWE-Bench Verified instances, compared to Full SWE-Bench and SWE-Repo tasks, showing a pattern of instance-specific memorization inside SWE-Bench repositories. Second, the sharp performance drop on external repos also persists, providing evidence for repository-bias memorization.
The persistence of these patterns in the filtered data is our most critical finding. It shows that the models’ ability to identify correct files is not due to simple text matching but is based on memorized knowledge, which includes instance-specific patterns and repository-biased patterns. This suggests that current benchmark scores are inflated by these memorization effects and don’t purely reflect a model’s transferable software engineering skills.
Model Match (%)
Claude 3.5 Sonnet 12.1
Claude 3.7 Sonnet 12.3
Claude 4 Sonnet 21.4
Claude 4 Opus 31.6
GPT-4.1 17.4
GPT-4o 24-05-13 18.4
GPT-4o 24-08-16 18.2
OpenAI o3-mini 11.7
Table 3: Instance-level verbatim match percentage across models on SWE-Bench Verified.
Figure 5: Comparison of models’ 5-gram overlap ratio against the ground truth code snippet on the function reproduction task across benchmarks.
@@ 4.3 Function Reproduction Result
We analyze model performance on the function reproduction task in this subsection. As shown in Figure 5, models achieve up to 34.9% 5-gram overlap ratio reproducing ground truth functions from SWE-Bench Verified compared to up to 18.2% on instances outside the SWE-Bench ecosystem111 Due to limited resources, we were only able to conduct experiments for o3-mini from among the OpenAI reasoning models. . With limited context provided, only function names and issue descriptions, models lack sufficient information to systematically derive correct implementations. High consecutive 5-gram overlap under these constraints requires reproducing exact textual sequences, indicating reliance on memorized training data rather than reasoning.
@@
: Performance reveals a clear hierarchy mirroring benchmark curation processes: SWE Verified (34.9%) >> SWE Full (28.7%) >> SWE Extra (18.2%). Critically, SWE Extra performance drops to the same level as external benchmarks (RefactorBench: 18.1%, Outside Tasks: 13.9%), despite originating from identical repositories as Verified. This pattern indicates that models have specifically memorized the curated, canonical solutions that human evaluators deemed high-quality, rather than developing general familiarity with repositories.
These results demonstrate that SWE-Bench Verified performance substantially reflects memorization of training sequences. The inability to reproduce functions from less-curated instances within the same repositories, combined with performance equivalence between SWE Extra and external benchmarks, reveals that apparent programming ability might result from an artifact of exposure to the specific, high-quality solutions selected for benchmark inclusion.
4.4 Direct Memorization on SWE-Bench Verified
Table 3 presents the results of our prefix completion experiment across eight SoTA language models, showing substantial evidence of training data contamination in SWE-Bench Verified, with the instance-level verbatim match percentage ranging from 11.7% to 31.6% across evaluated models.
The Claude family exhibits a monotonic increase in memorization rates corresponding to model generation, rising from 12.1% to 31.6%. This progression suggests a strong correlation between model capacity and memorization. Notably, Claude 4 Opus demonstrates verbatim generation for nearly a third of all test instances. In contrast, the GPT family displays relatively stable memorization rates (17.4%-18.4%), indicating potential differences in training data curation or architectural choices that affect memorization. OpenAI o3-mini exhibits the lowest repro rate at 11.7%.
These results have significant implications for benchmark validity. The ability of models to reproduce exact code sequences when provided merely with contextual prefixes, without any problem description or bug identification, strongly suggests that performance on these instances reflects memorization rather than algorithmic reasoning or program comprehension. This finding strengthens our hypothesis that a substantial portion of reported performance gains on SWE-Bench may be attributed to memorization rather than genuine advances in programming capabilities.
@@ 5.1 Coding Benchmarks and Agents
Various coding benchmarks have been proposed to evaluate LLMs and LLM agents’, including SWE-Bench, BigCodeBench, SWE-Gym, EvoEval (Jimenez et al. 2024; Zhuo et al. 2024; Xia, Deng, and Zhang 2024). Among them, SWE-Bench has emerged as the most popular evaluation method for LLM’s coding abilities (Jimenez et al. 2024). The benchmark consists of real GitHub issues and their corresponding patches, providing a seemingly realistic evaluation setting. However, the public nature of GitHub data raises concerns about potential exposure during training. Recent agent systems have shown dramatic improvements on SWE-Bench (Yang et al. 2024; Wang et al. 2024), but the extent to which these improvements reflect genuine problem-solving versus pattern matching remains unclear.
@@ 5.2 Benchmark Contamination
The issue of benchmark contamination, where models are inadvertently trained on eval data, is a well-documented problem that can lead to inflated performance metrics without genuine understanding (Zhou et al. 2023). To combat this in the coding domain, prior work has developed two primary strategies: task mutation and metric-based probing.
Task mutation approaches aim to detect memorization by altering existing benchmarks. For example, EvoEval (Xia, Deng, and Zhang 2024) uses LLMs to create semantic-altering transformations of HumanEval (Chen et al. 2021) problems, while other methods (Chen et al. 2025) use AST-level mutations to measure how performance changes with structural code changes. Similarly, TaskEval (Tambon et al. 2025) measures prompt sensitivity by creating variations of a task. These methods are effective at revealing overfitting on specific problem formulations.
Metric-based probing approaches, in contrast, analyze model outputs without changing the task itself. These techniques (Li 2023; Ramos et al. 2025) include using perplexity or other statistical measures like Negative Log Likelihood (NLL) and n-gram accuracy to infer if model has likely seen a specific bug or solution before. However, these measures cannot be applied to commercial models as we do not have access to the model’s hidden states. N-gram similarity, while applicable, is a noisy indicator for complex code patches; any functionally correct solution will inevitably share text overlap with ground truth, making it hard to distinguish genuine problem-solving from verbatim memorization.
Our work introduces a novel cross benchmark analysis framework that addresses these limitations by comparing models across related benchmarks as memorization proxies. Rather than requiring knowledge of training data exposure, we systematically compare model performance across benchmark variants and external evaluations to detect suspicious performance patterns. This provides a systematic, reusable framework for detecting memorization in any coding benchmark without requiring access to training data or model internals.